A NUMBER DECREASED BY 1/3 OF THE NUMBER IS 36. FIND THE NUMBER.

X - 1/3 = 36

DO I ADD?

Answer by longjonsilver(2297)

let number be x
So that a third of the number is x/3

a number is decreased by a third of the number:
x - x/3 and this equals 36

x - x/3 = 36
2x/3 = 36
x/3 = 18
x = 54

jon.